# Normal coactions extend to the C*-envelope

### 13 Mar 2024 v2

@@ 3. Existence of coactions on C*-envelopes
Suppose $\mathcal{A}$ has a contractive approximate identity, and assume that $\delta\colon G\downtouparrow\mathcal{A}$ is a coaction by a discrete group $G$. Then, the coaction crossed product $\mathcal{A}\rtimes_{\delta}G$ has a contractive approximate identity.
@@
First, suppose we have a contractive approximate identity $\{a_{\alpha}\}_{\alpha\in I}$ of $\mathcal{A}$ and let $\mathbb{E}_{e}$ be the canonical conditional expectation onto $\mathcal{A}_{e}$, as described in § 2.1. We will show that $\{\mathbb{E}_{e}(a_{\alpha})\}_{\alpha\in I}$, which is a net in $\mathcal{A}_{e}$, is a contractive approximate identity for $\mathcal{A}$. It is clearly contractive, because $\mathbb{E}_{e}$ is a contractive idempotent. Now, for $a\in\mathcal{A}_{g}$ and $g\in G$ we first show that $a\mathbb{E}_{e}(a_{\alpha})\rightarrow a$. Indeed,
@@
Hence, we may assume without loss of generality that $\{a_{\alpha}\}_{\alpha\in I}$ is a contractive approximate unit for $\mathcal{A}$ in $\mathcal{A}_{e}$. Let $\{\chi_{F}\}_{F}$ be the net of characteristic functions supported on finite subsets $F\subseteq G$, which we denote by $\text{Fin}(G)$. We will show that the net $\{a_{\alpha}\chi_{F}\}_{(\alpha,F)\in I\times\text{Fin}(G)}$ is a contractive approximate identity for $\mathcal{A}\rtimes_{\delta}G$.
It suffices to show that $\{a_{\alpha}\chi_{F}\}_{(\alpha,F)\in I\times\text{Fin}(G)}$ is an approximate identity on generators of the form $\delta_{\lambda}(a)j_{c_{0}(G)}(\chi_{E})$, where $a\in\mathcal{A}_{g}$ for $g\in G$, and $E\subseteq G$ is finite. In this case, if we take $F$ finite such that $gE,E\subseteq F$, and by multiplying on the right we see that
@@
Let $\mathcal{A}$ be an operator algebra with a contractive approximate identity, and let $\kappa_{\mathcal{A}}\colon\mathcal{A}\to C^{*}_{\textup{env}}(\mathcal{A})$ be the canonical completely isometric inclusion. Suppose $\delta\colon G\downtouparrow\mathcal{A}$ is a normal coaction of a discrete group $G$. Then, there exists a normal coaction $\delta_{\textup{env}}\colon G\downtouparrow C_{\textup{env}}^{*}(\mathcal{A})$ such that $\delta_{\textup{env}}\circ\kappa_{\mathcal{A}}=(\kappa_{\mathcal{A}}\otimes\textup{id}_{C^{*}(G)})\circ\delta$.
@@
Let $\Psi\colon\mathcal{A}\rtimes_{\delta}G\rtimes^{r}_{\hat{\delta}}G\to\delta_{\lambda}(\mathcal{A})\otimes\mathbb{K}$ be the Katayama isomorphism from Theorem 3.3, and let $\Psi^{-1}_{\textup{env}}\colon C_{\textup{env}}^{*}(\delta_{\lambda}(\mathcal{A})\otimes\mathbb{K})\to C_{\textup{env}}^{*}(\mathcal{A}\rtimes_{\delta}G\rtimes^{r}_{\hat{\delta}}G)$ be the *-isomorphism induced from $\Psi^{-1}$ between the C*-envelopes. By the co-universal property of the C*-envelope, the dual action $\hat{\delta}$ of $G$ on $\mathcal{A}\rtimes_{\delta}G$ extends to an action, also denoted $\hat{\delta}$, of $G$ on $C_{\textup{env}}^{*}(\mathcal{A}\rtimes_{\delta}G)$ (see [KR19, Lemma 3.4]). By Lemma 3.4, $\mathcal{A}\rtimes_{\delta}G$ has a contractive approximate identity, so [Ka17, Theorem 2.5] gives us a *-isomorphism $\theta\colon C^{*}_{\textup{env}}(\mathcal{A}\rtimes_{\delta}G\rtimes^{r}_{\hat{\delta}}G)\to C^{*}_{\textup{env}}(\mathcal{A}\rtimes_{\delta}G)\rtimes^{r}_{\hat{\delta}}G$ that maps generators to generators. We summarize the above discussion in the following commutative diagram.
@@
Let $\mathcal{A}$ be an operator algebra with a contractive approximate identity, and suppose $\delta\colon G\downtouparrow\mathcal{A}$ is a normal coaction of a discrete group $G$. Then, the canonical map $C_{\textup{env}}^{*}(\mathcal{A},\delta)\to C_{\textup{env}}^{*}(\mathcal{A})$ is a *-isomorphism.
@@ 4.1. The boundary quotient is a C*-cover
$[X,\chi]\coloneq\{[c,\chi]:c\in X,\chi(\mathfrak{d}(c)\mathfrak{C})=1\}\subseteq(I_{l}\ltimes\partial\Omega)_{\chi}.$
Since $\mathfrak{C}$ is cancellative, we have $[c,\chi]=[d,\chi]$ if and only if $c=d$. Indeed, the equality $[c,\chi]=[d,\chi]$ implies that there exists $Y\in\chi^{-1}(1)$ such that $cy=dy$ for $y\in Y$ and, by right cancellation, this implies that $c=d$.
@@
where $\textup{fix}(s)\coloneq\{x\in\textup{dom}(s):s(x)=x\}$. On the other hand, we have that $E_{r}(1_{[s,\Omega(\textup{dom}(s))]}))=1_{[s,\Omega(\textup{dom}(s))]\cap\Omega}$. By [EP16, Proposition 3.14], we get $[s,\Omega(\textup{dom}(s))]\cap\Omega=\mathcal{F}_{s}$ where we define $\mathcal{F}_{s}\coloneq\bigcup_{X\in\mathcal{J},X\subseteq\textup{fix}(s)}\Omega(X)$. Since $I_{l}\ltimes\Omega$ is Hausdorff, [Li23, Lemma 4.1(i)] implies that there exists a finite (possibly empty) subset $F\subseteq\mathcal{J}^{\times}$ such that $\textup{fix}(s)=\bigcup_{X\in F}X$, so that $\mathcal{F}_{s}=\bigcup_{X\in F}\Omega(X)$. By the inclusion-exclusion principle, we have that
@@
Since both $E_{r}$ and $\prod_{\chi}\Phi_{\chi}$ are faithful, in order to prove $\vartheta$ is injective, it will suffice to prove that $\vartheta|_{C_{0}(\Omega)}$ is injective. By [CEL15, Proposition 5.6.21], the kernel of $\vartheta|_{C_{0}(\Omega)}$ is generated by the projections $1_{\Omega(X)}-1_{\bigcup_{i=1}^{n}\Omega(X_{i})}$, where $X,X_{1},\ldots,X_{n}\in\mathcal{J}$ with $X_{i}\subseteq X$ are such that $[X,\chi]=\bigcup_{i=1}^{n}[X_{i},\chi]$ for all $\chi\in\partial\Omega$. Thus, it will suffice to show that $\vartheta_{\chi}$ is nonzero on all such projections. Fix such a projection $1_{\Omega(X)}-1_{\bigcup_{i=1}^{n}\Omega(X_{i})}$ and note that it is equal to zero if and only if $X=\bigcup_{i=1}^{n}X_{i}$. So we need only prove that $X=\bigcup_{i=1}^{n}X_{i}$ whenever we have $[X,\chi]=\bigcup_{i=1}^{n}[X_{i},\chi]$ for every $\chi\in\partial\Omega$. Take $x\in X$, and choose $\chi\in\Omega_{\textup{max}}$ such that $\chi(\mathfrak{d}(x)\mathfrak{C})=1$ (such a $\chi$ exists by [Li23, Lemma 2.21(ii)]). Then, since $[x,\chi]\in[X,\chi]$ and we have the equality $[X,\chi]=\bigcup_{i=1}^{n}[X_{i},\chi]$, we get that $[x,\chi]\in[X_{i},\chi]$ for some $i$, so $x\in X_{i}$ by right cancellation. We conclude that $\vartheta|_{C_{0}(\Omega)}$ is injective. ∎
@@
Since $(I_{l}\ltimes\Omega)_{\chi}=(I_{l}\ltimes\partial\Omega)_{\chi}$ for all $\chi\in\partial\Omega$, we have $q_{\partial}=\oplus_{\chi\in\partial\Omega}\rho_{\chi}$. For each $\chi\in\partial\Omega$, let $P_{\chi}\colon\ell^{2}((I_{l}\ltimes\Omega)_{\chi}))\to\ell^{2}([\mathfrak{C},\chi])$ be the orthogonal projection associated with the inclusion $[\mathfrak{C},\chi]\subseteq(I_{l}\ltimes\Omega)_{\chi}$. A straightforward calculation shows that
@@
for all $c\in\mathfrak{C}$ and $[d,\chi]\in[\mathfrak{C},\chi]$. Hence, the subspace $\ell^{2}([\mathfrak{C},\chi])$ is invariant under $\rho_{\chi}(1_{[c,\Omega(\mathfrak{d}(c)\mathfrak{C})]})$, and $P_{\chi}\rho_{\chi}(1_{[c,\Omega(\mathfrak{d}(c)\mathfrak{C})]})P_{\chi}=\vartheta_{\chi}(1_{[c,\Omega(\mathfrak{d}(c)\mathfrak{C})]})$. It follows that $\bigoplus_{\chi\in\partial\Omega}\ell^{2}([\mathfrak{C},\chi])$ is invariant under $\mathcal{A}_{r}(\mathfrak{C})$, so we obtain an algebra homomorphism $\psi\colon\mathcal{A}_{r}(\mathfrak{C})\to\mathbb{B}\left(\textstyle{\bigoplus}_{\chi\in\partial\Omega}\ell^{2}([\mathfrak{C},\chi])\right)$ given by
@@ 4.2. From functors to coactions
such that $\kappa_{\rho}([s,\chi])=\bar{\rho}(s)$ for all $[s,\chi]\in I_{l}\ltimes\Omega$.
@@
If $[s,\chi]=[t,\chi]$ in $I_{l}\ltimes\Omega$, then there exists $X\in\chi^{-1}(1)$ with $s\circ\textup{id}_{X}=t\circ\textup{id}_{X}\neq 0$, so that $\bar{\rho}(s)=\bar{\rho}(s)\bar{\rho}(\textup{id}_{X})=\bar{\rho}(s\circ\textup{id}_{X})=\bar{\rho}(t\circ\textup{id}_{X})=\bar{\rho}(t)$. Thus, there is a well-defined map $\kappa_{\rho}\colon I_{l}\ltimes\Omega\to\mathcal{U}(\mathfrak{G})$ such that $\kappa_{\rho}([s,\chi])=\bar{\rho}(s)$ for all $[s,\chi]\in I_{l}\ltimes\Omega$. If $\chi\in\Omega$ and $t,s\in I_{l}$ with $[t,s.\chi][s,\chi]=[ts,\chi]$, then $ts\neq 0$ and
@@
[…]
@@
Given a category $\mathfrak{C}$, we let $\textup{Env}(\mathfrak{C})$ be the enveloping groupoid of $\mathfrak{C}$ in the sense of [Deh15, Definition II.3.3] (note that $\textup{Env}(\mathfrak{C})$ is called the fundamental groupoid in [BKQ21]), and let $\rho_{u}\colon\mathfrak{C}\to\textup{Env}(\mathfrak{C})$ be the canonical functor.
@@
[…]
Every element of $I_{l}\ltimes\partial\Omega$ can be written as $[cd^{-1},\chi]$ for some $c,d\in\mathfrak{C}$ with $\mathfrak{d}(c)=\mathfrak{d}(d)$ and $\chi\in\partial\Omega$ with $\chi(d\mathfrak{C})=1$ (see, e.g., [Li23, Lemma 2.19]). Take such an element $[cd^{-1},\chi]$ and suppose $\partial\kappa([cd^{-1},\chi])=e$. This means that $\rho_{u}(c)=\rho_{u}(d)$, so by commutativity of (LABEL:diag:deg), we have $\deg(c)=\deg(d)$. Assuming that $[cd^{-1},\chi]$ is isotropy, i.e., that $cd^{-1}.\chi=\chi$, we have $1=\chi(d\mathfrak{C})=cd^{-1}.\chi(d\mathfrak{C})=\chi(dc^{-1}(c\mathfrak{C}\cap d\mathfrak{C}))$. In particular, $c\mathfrak{C}\cap d\mathfrak{C}\neq\emptyset$, and [OP20, Lemma 8.2] implies that $c=d$. Therefore, $[cd^{-1},\chi]$ is a unit, and this shows that $\ker(\partial\kappa)$ is principal. ∎

### 17 Mar 2026 v3

@@ 3. Existence of coactions on C*-envelopes
Suppose $\mathcal{A}$ has a contractive self-adjoint approximate identity, and assume that $\delta\colon G\downtouparrow\mathcal{A}$ is a coaction by a discrete group $G$. Then, the coaction crossed product $\mathcal{A}\rtimes_{\delta}G$ has a contractive self-adjoint approximate identity.
@@
First, suppose we have a contractive self-adjoint approximate identity $\{a_{\alpha}\}_{\alpha\in I}$ of $\mathcal{A}$ and let $\mathbb{E}_{e}$ be the canonical conditional expectation onto $\mathcal{A}_{e}$, as described in § 2.1. We will show that $\{\mathbb{E}_{e}(a_{\alpha})\}_{\alpha\in I}$, which is a net in $\mathcal{A}_{e}$, is a contractive self-adjoint approximate identity for $\mathcal{A}$. It is clearly contractive and self-adjoint, because $\mathbb{E}_{e}$ is a contractive idempotent. Now, for $a\in\mathcal{A}_{g}$ and $g\in G$ we first show that $a\mathbb{E}_{e}(a_{\alpha})\rightarrow a$. Indeed,
@@
Hence, we may assume without loss of generality that $\{a_{\alpha}\}_{\alpha\in I}$ is a contractive self-adjoint approximate unit for $\mathcal{A}$ which is in $\mathcal{A}_{e}$. Let $\{\chi_{F}\}_{F}$ be the net of characteristic functions supported on finite subsets $F\subseteq G$, which we denote by $\text{Fin}(G)$. We will show that the net of real parts $\{\frac{\chi_{F}a_{\alpha}+a_{\alpha}\chi_{F}}{2}\}_{(\alpha,F)\in I\times\text{Fin}(G)}$, which is comprised of contractive self-adjoint elements, is an approximate identity for $\mathcal{A}\rtimes_{\delta}G$.
It suffices to show that $\{a_{\alpha}\chi_{F}\}_{(\alpha,F)\in I\times\text{Fin}(G)}$ is both a left and a right approximate identity on generators of the form $\delta_{\lambda}(a)j_{c_{0}(G)}(\chi_{E})$, where $a\in\mathcal{A}_{g}$ for $g\in G$, and $E\subseteq G$ is finite, and the proof for $\{\chi_{F}a_{\alpha}\}$ is then handled similarly. So, for $\{a_{\alpha}\chi_{F}\}_{(\alpha,F)\in I\times\text{Fin}(G)}$, if we take $F$ finite such that $gE,E\subseteq F$, and by multiplying on the right we see that
@@
Let $\mathcal{A}$ be an operator algebra with a contractive self-adjoint approximate identity, and let $\kappa_{\mathcal{A}}\colon\mathcal{A}\to C^{*}_{\textup{env}}(\mathcal{A})$ be the canonical completely isometric inclusion. Suppose $\delta\colon G\downtouparrow\mathcal{A}$ is a normal coaction of a discrete group $G$. Then, there exists a normal coaction $\delta_{\textup{env}}\colon G\downtouparrow C_{\textup{env}}^{*}(\mathcal{A})$ such that $\delta_{\textup{env}}\circ\kappa_{\mathcal{A}}=(\kappa_{\mathcal{A}}\otimes\textup{id}_{C^{*}(G)})\circ\delta$.
@@
Let $\Psi\colon\mathcal{A}\rtimes_{\delta}G\rtimes^{r}_{\hat{\delta}}G\to\delta_{\lambda}(\mathcal{A})\otimes\mathbb{K}$ be the Katayama isomorphism from Theorem 3.3, and let $\Psi^{-1}_{\textup{env}}\colon C_{\textup{env}}^{*}(\delta_{\lambda}(\mathcal{A})\otimes\mathbb{K})\to C_{\textup{env}}^{*}(\mathcal{A}\rtimes_{\delta}G\rtimes^{r}_{\hat{\delta}}G)$ be the *-isomorphism induced from $\Psi^{-1}$ between the C*-envelopes. By the co-universal property of the C*-envelope, the dual action $\hat{\delta}$ of $G$ on $\mathcal{A}\rtimes_{\delta}G$ extends to an action, also denoted $\hat{\delta}$, of $G$ on $C_{\textup{env}}^{*}(\mathcal{A}\rtimes_{\delta}G)$ (see [KR19, Lemma 3.4]). By Lemma 3.4, $\mathcal{A}\rtimes_{\delta}G$ has a contractive self-adjoint approximate identity, so [Ka17, Theorem 2.5] gives us a *-isomorphism $\theta\colon C^{*}_{\textup{env}}(\mathcal{A}\rtimes_{\delta}G\rtimes^{r}_{\hat{\delta}}G)\to C^{*}_{\textup{env}}(\mathcal{A}\rtimes_{\delta}G)\rtimes^{r}_{\hat{\delta}}G$ that maps generators to generators. We summarize the above discussion in the following commutative diagram.
@@
Let $\mathcal{A}$ be an operator algebra with a contractive self-adjoint approximate identity, and suppose $\delta\colon G\downtouparrow\mathcal{A}$ is a normal coaction of a discrete group $G$. Then, the canonical map $C_{\textup{env}}^{*}(\mathcal{A},\delta)\to C_{\textup{env}}^{*}(\mathcal{A})$ is a *-isomorphism.
Remark 3.7.
Although we do not discuss product systems of C*-correspondences over group-embeddable monoids in this paper, it is worthwhile to sketch the simplified proof of [Seh22, Theorem 5.1] that we may now obtain. More precisely, we can now dispense with the context-specific techniques used in [Seh22, Lemma 3.7] and [Seh22, Section 4] by appealing to Theorem 3.5 and Corollary 3.6 combined with [Seh22, Corollary 3.5].
To provide a sketch of the proof, in what follows we adhere to terminology and conventions from [Seh22]. Let $\mathcal{E}$ be a product system of C*-correspondences over a monoid $P$ embeddable in a group $G$. To show that the C*-envelope of the tensor algebra $\mathcal{T}_{\lambda}^{+}(\mathcal{E})$ is the reduced Fell bundle C*-algebra completion of $([A\times_{\mathcal{E}}P]_{g})_{g\in G}$ (which is denoted by $C^{*}_{r}(([A\times_{\mathcal{E}}P]_{g})_{g\in G})$) we first note that by Theorem 3.5 we have that $C^{*}_{\textup{env}}(\mathcal{T}_{\lambda}^{+}(\mathcal{E}))$ carries a canonical normal coaction by $G$ induced by the natural one on $\mathcal{T}_{\lambda}^{+}(\mathcal{E})$. By [Seh22, Corollary 3.5] the canonical quotient $\mathcal{T}_{\lambda}(\mathcal{E})\rightarrow C^{*}_{r}(([A\times_{\mathcal{E}}P]_{g})_{g\in G})$ is completely isometric on $\mathcal{T}_{\lambda}^{+}(\mathcal{E})$, so that by the co-universal property of $C^{*}_{\textup{env}}(\mathcal{T}_{\lambda}^{+}(\mathcal{E}))$ we have a canonical quotient map $C^{*}_{r}(([A\times_{\mathcal{E}}P]_{g})_{g\in G})\rightarrow C^{*}_{\textup{env}}(\mathcal{T}_{\lambda}^{+}(\mathcal{E}))$. But now, since this quotient map is automatically equivariant with respect to the canonical normal coactions on these C*-algebras, we get that this quotient map intertwines the natural faithful conditional expectations on these two C*-algebras arising from the respective normal coactions. A standard conditional expectation argument together with [Seh18, Theorem 3.10] then shows that this canonical quotient map is injective. Corollary 3.6 then furnishes a complete proof of [Seh22, Theorem 5.1].
@@ 4.1. The boundary quotient is a C*-cover
$X,\chi]\coloneq\{[c,\chi]:c\in X,\chi(\mathfrak{d}(c)\mathfrak{C})=1\}\subseteq(I_{l}\ltimes\partial\Omega)_{\chi}.$
Since $\mathfrak{C}$ is cancellative, we have $c,\chi]=[d,\chi]$ if and only if $c=d$. Indeed, the equality $c,\chi]=[d,\chi]$ implies that there exists $Y\in\chi^{-1}(1)$ such that $cy=dy$ for $y\in Y$ and, by right cancellation, this implies that $c=d$.
@@
where $\textup{fix}(s)\coloneq\{x\in\textup{dom}(s):s(x)=x\}$. On the other hand, we have that $E_{r}(1_{[s,\Omega(\textup{dom}(s))]}))=1_{[s,\Omega(\textup{dom}(s))]\cap\Omega}$. By [EP16, Proposition 3.14], we get $s,\Omega(\textup{dom}(s))]\cap\Omega=\mathcal{F}_{s}$ where we define $\mathcal{F}_{s}\coloneq\bigcup_{X\in\mathcal{J},X\subseteq\textup{fix}(s)}\Omega(X)$. Since $I_{l}\ltimes\Omega$ is Hausdorff, [Li23, Lemma 4.1(i)] implies that there exists a finite (possibly empty) subset $F\subseteq\mathcal{J}^{\times}$ such that $\textup{fix}(s)=\bigcup_{X\in F}X$, so that $\mathcal{F}_{s}=\bigcup_{X\in F}\Omega(X)$. By the inclusion-exclusion principle, we have that
@@
Since both $E_{r}$ and $\prod_{\chi}\Phi_{\chi}$ are faithful, in order to prove $\vartheta$ is injective, it will suffice to prove that $\vartheta|_{C_{0}(\Omega)}$ is injective. By [CEL15, Proposition 5.6.21], the kernel of $\vartheta|_{C_{0}(\Omega)}$ is generated by the projections $1_{\Omega(X)}-1_{\bigcup_{i=1}^{n}\Omega(X_{i})}$, where $X,X_{1},\ldots,X_{n}\in\mathcal{J}$ with $X_{i}\subseteq X$ are such that $X,\chi]=\bigcup_{i=1}^{n}[X_{i},\chi]$ for all $\chi\in\partial\Omega$. Thus, it will suffice to show that $\vartheta_{\chi}$ is nonzero on all such projections. Fix such a projection $1_{\Omega(X)}-1_{\bigcup_{i=1}^{n}\Omega(X_{i})}$ and note that it is equal to zero if and only if $X=\bigcup_{i=1}^{n}X_{i}$. So we need only prove that $X=\bigcup_{i=1}^{n}X_{i}$ whenever we have $X,\chi]=\bigcup_{i=1}^{n}[X_{i},\chi]$ for every $\chi\in\partial\Omega$. Take $x\in X$, and choose $\chi\in\Omega_{\textup{max}}$ such that $\chi(\mathfrak{d}(x)\mathfrak{C})=1$ (such a $\chi$ exists by [Li23, Lemma 2.21(ii)]). Then, since $x,\chi]\in[X,\chi]$ and we have the equality $X,\chi]=\bigcup_{i=1}^{n}[X_{i},\chi]$, we get that $x,\chi]\in[X_{i},\chi]$ for some $i$, so $x\in X_{i}$ by right cancellation. We conclude that $\vartheta|_{C_{0}(\Omega)}$ is injective. ∎
@@
Since $(I_{l}\ltimes\Omega)_{\chi}=(I_{l}\ltimes\partial\Omega)_{\chi}$ for all $\chi\in\partial\Omega$, we have $q_{\partial}=\oplus_{\chi\in\partial\Omega}\rho_{\chi}$. For each $\chi\in\partial\Omega$, let $P_{\chi}\colon\ell^{2}((I_{l}\ltimes\Omega)_{\chi}))\to\ell^{2}([\mathfrak{C},\chi])$ be the orthogonal projection associated with the inclusion $\mathfrak{C},\chi]\subseteq(I_{l}\ltimes\Omega)_{\chi}$. A straightforward calculation shows that
@@
for all $c\in\mathfrak{C}$ and $d,\chi]\in[\mathfrak{C},\chi]$. Hence, the subspace $\ell^{2}([\mathfrak{C},\chi])$ is invariant under $\rho_{\chi}(1_{[c,\Omega(\mathfrak{d}(c)\mathfrak{C})]})$, and $P_{\chi}\rho_{\chi}(1_{[c,\Omega(\mathfrak{d}(c)\mathfrak{C})]})P_{\chi}=\vartheta_{\chi}(1_{[c,\Omega(\mathfrak{d}(c)\mathfrak{C})]})$. It follows that $\bigoplus_{\chi\in\partial\Omega}\ell^{2}([\mathfrak{C},\chi])$ is invariant under $\mathcal{A}_{r}(\mathfrak{C})$, so we obtain an algebra homomorphism $\psi\colon\mathcal{A}_{r}(\mathfrak{C})\to\mathbb{B}\left(\textstyle{\bigoplus}_{\chi\in\partial\Omega}\ell^{2}([\mathfrak{C},\chi])\right)$ given by
@@ 4.2. From functors to coactions
such that $\kappa_{\rho}([s,\chi])=\bar{\rho}(s)$ for all $s,\chi]\in I_{l}\ltimes\Omega$.
@@
If $s,\chi]=[t,\chi]$ in $I_{l}\ltimes\Omega$, then there exists $X\in\chi^{-1}(1)$ with $s\circ\textup{id}_{X}=t\circ\textup{id}_{X}\neq 0$, so that $\bar{\rho}(s)=\bar{\rho}(s)\bar{\rho}(\textup{id}_{X})=\bar{\rho}(s\circ\textup{id}_{X})=\bar{\rho}(t\circ\textup{id}_{X})=\bar{\rho}(t)$. Thus, there is a well-defined map $\kappa_{\rho}\colon I_{l}\ltimes\Omega\to\mathcal{U}(\mathfrak{G})$ such that $\kappa_{\rho}([s,\chi])=\bar{\rho}(s)$ for all $s,\chi]\in I_{l}\ltimes\Omega$. If $\chi\in\Omega$ and $t,s\in I_{l}$ with $t,s.\chi][s,\chi]=[ts,\chi]$, then $ts\neq 0$ and
@@
$C_{r}^{*}(I_{l}\ltimes\partial\Omega)$$C_{\textup{env}}^{*}(\mathcal{A}_{r}(\mathfrak{C}))$$C_{r}^{*}(\ker(\partial\kappa_{\rho}))$$C_{\textup{env}}^{*}(\mathcal{A}_{r}(\mathfrak{C}))_{e}\makebox[0.0pt][l]{\,,}$$\scriptstyle{\partial E}$$\scriptstyle{\pi_{\textup{env}}}$$\scriptstyle{\Psi}$$\scriptstyle{\pi_{\textup{env}}|_{C_{r}^{*}(\ker(\partial\kappa_{\rho}))}}$
@@
Given a category $\mathfrak{C}$, we let $\textup{Env}(\mathfrak{C})$ be the enveloping groupoid of $\mathfrak{C}$ in the sense of [Deh15, Definition II.3.3], and let $\rho_{u}\colon\mathfrak{C}\to\textup{Env}(\mathfrak{C})$ be the canonical functor (note that $\textup{Env}(\mathfrak{C})$ is called the fundamental groupoid in [BKQ21]). The pair $(\textup{Env}(\mathfrak{C}),\rho_{u})$ is uniquely determined up to isomorphism by the following universal property: for every functor $F\colon\mathfrak{C}\to\mathcal{H}$, where $\mathcal{H}$ is a groupoid, there exists a unique functor $\widetilde{F}\colon\textup{Env}(\mathfrak{C})\to\mathcal{H}$ such that $F=\rho_{u}\circ\widetilde{F}$.
@@
(4.6) $\mathfrak{C}$$\textup{Env}(\mathfrak{C})$$G\makebox[0.0pt][l]{\,.}$$\scriptstyle{\deg}$$\scriptstyle{\rho_{u}}$$\scriptstyle{\rho_{u}^{\prime}}$
Every element of $I_{l}\ltimes\partial\Omega$ can be written as $cd^{-1},\chi]$ for some $c,d\in\mathfrak{C}$ with $\mathfrak{d}(c)=\mathfrak{d}(d)$ and $\chi\in\partial\Omega$ with $\chi(d\mathfrak{C})=1$ (see, e.g., [OP20, Lemma 4.8]). Take such an element $cd^{-1},\chi]$ and suppose $\partial\kappa([cd^{-1},\chi])=e$. This means that $\rho_{u}(c)=\rho_{u}(d)$, so by commutativity of (4.6), we have $\deg(c)=\deg(d)$. Assuming that $cd^{-1},\chi]$ is isotropy, i.e., that $cd^{-1}.\chi=\chi$, we have $1=\chi(d\mathfrak{C})=cd^{-1}.\chi(d\mathfrak{C})=\chi(dc^{-1}(c\mathfrak{C}\cap d\mathfrak{C}))$. In particular, $c\mathfrak{C}\cap d\mathfrak{C}\neq\emptyset$, and [OP20, Lemma 8.2] implies that $c=d$. Therefore, $cd^{-1},\chi]$ is a unit, and this shows that $\ker(\partial\kappa)$ is principal. ∎